An object is shoved off a table at 3.0 m/s and hits the floor 0.75m below the table top. What is its speed just before it hits the floor?
Answer #1

t = t= 0.391s égt² =) (ots) = { (-9.8) t - ta 3 mis = gt = (-9.8) (0.391) = Vot - Vo -3.83 mis & sortu A .. speed = = 32 +(3.

Kindly upvote:)